Fermi National Laboratory


Accelerator Update

Friday May 28
The day shift began with the Tevatron (TeV) in cooldown and recovery, the Antiproton Source (Pbar) stacking, with the Recycler conducting studies, and with MiniBooNE taking beam.

By 8:34 AM, Linac technicians had repaired problems with two RF stations, (LRF2 and KRF2).

Booster suffered a vacuum burst at 8:55 AM. Everything reset okay.

LRF2 tripped off at 9:32 AM and then again 30 minutes later.

A power supply tech repaired a Meson MCenter quadrupole power supply at 9:43 AM.

Recycler experts completed their transfers of antiprotons to the Main Injector (MI) at 10:36 AM. Operators prepared MI for an access.

An expert investigated why the Booster chopper was occasionally missing pulses and sparking at 11:23 AM. He replaced a trigger card.

MI experts reported at noon that they found a leak on the MIRF8 cavity. They completed their repairs about two hours later.

At 12:14 PM, operations put current on to the TeV buss.

At 2:03 PM, the TeV suffered a sector B0 quench. Experts investigated.

Operations turned on MI at 2:48 PM, established beam to MiniBooNE at 2:52 PM, and established beam to Pbar at 2:56 PM.

At 3:06 PM the Booster chopper started missing pulse again. The expert reported at 4:05 PM that he had repaired the chopper.

Operations fired up the TeV and at 7:15 PM put the TeV through a dry squeeze followed by a wet squeeze and then TeV orbit checks.

Saturday May 29
The midnight shift began with Pbar stacking, with TeV experts conducting studies, and with MiniBooNE taking beam.

At 12:13 AM, three Booster beam valves closed, a RF station, the chopper, and a power supply tripped off. Everything reset okay.

Operations loaded a proton-only study-store at 1:19 AM.

Around 2 AM, the TeV aborted. There was no beam in the machine. Operators investigated, but could find no reason for the abort. An expert advised the operators to change out some timing modules.

Operations put the TeV through another dry squeeze at 3:18 AM.

Operations put the TeV into shot setup at 3:42 AM.

At 3:58 AM, a TeV Lambertson (LAM52) began tripping off on excessive current ripple. Operations tried adjusting the 2D timing event on the Time Line Generator (TLG), but it didn't help. Experts came in and investigated.

At 4:30 AM, a Booster beam valve closed and the chopper tripped off. Operators discovered that a vacuum controller had tripped off and wouldn't reset. An expert came in to investigate.

At 5:21 AM, Operations decided that a TeV sextupole power supply (SEXTHPS) had been tripping off too often and called techs to make repairs.

Operations reestablished beam to MiniBooNE at 6:27 AM.

A TeV power supply tech said that they had replaced a regulator chassis for LAM52. At 12:42 PM, the tech said it was ready to go. Operations put the TeV through a dry squeeze at 12:44 PM.

The TeV tune-up was halted at 1:34 PM due to a beam abort. A TeV expert thinks a collimator in sector D1 may be skewed and stuck due to a limit switch problem. An expert came in.

MI and Pbar experts completed the machine tune-up at 1:55 PM.

At 2:18 PM, SEXHPS tripped off for the first of four times in two hours. Operations left the power supply off after the last trip and called experts in.

Operations prepared the TEV for an access and at 4:38 PM found that a collimator had definitely gone out of alignment.

If you'll notice the positions of the up and downstream bellows you can see that the collimator was in the beam.

Two Booster RF stations (BRF9 & 10) tripped off at 4:57 PM due to high pressure. They both reset okay.

At 5:08 PM, two power supply experts began working on SEXHPS again.

At 5:38 PM, D-zero (D0) experimenters discovered a problem with the calorimeter that would keep them from taking data. Operators prepared their collision hall for an access at 6 PM. They were out of access by 6:35 PM.

Power supply techs completed their work on SEXHPS at 6:45 PM.

Operations put the TeV through a dry squeeze at 6:52 PM and through a wet squeeze at 7:26 PM.

Operations put the TeV into shot setup around 8:30 PM.

Operations reported at 10:07 PM that the TeV sector A4 cryo system was having problems. CHL cryo system techs called in an expert to investigate.

Cryo reported at 11:03 PM that they would have to replace a reciprocating compressor, which was a four to six hour job.

At 11:11 PM, Operations reported that they lost 120mA of the stack while trying to return to the stacking lattice due to an emittance problem.

At 11:39 PM, the TeV suffered a sector F2 & 3 quench.

Sunday May 30
The midnight shift began with TeV recovering from a quench, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

Operations reported that the TeV cryo systems at sectors F2 & 3 had recovered by 12:45 AM.

Operations reported that the TeV turn on began at 3:54 AM.

Operations put the TeV into shot setup at 8:31 AM. All Switchyard (SY) beam is turned off while the TeV is in shot setup.

Operations put the shot setup on hold while a severe storm approached at 9:40 AM. Shot setup resumed at 11:13 AM.

At 11:55 AM, Operations reported that they were having Pbar to MI beamline transfer problems.

Operations established store 3547 at 12:13 PM with an initial luminosity of 25.01E30.

Pbar resumed stacking at 1:10 PM.

At 1:35 PM, a Booster RF station (BRF18) tripped of for the first of three times in two hours.

At 2:16 PM, a Pbar Debuncher RF station (DRF1) developed a driver fault and wouldn't reset. After talking with an expert, operators replaced a power supply. The station was operational by 4:47 PM.

A Booster RF station (BRF16) tripped off for the first of four times in about thirty minutes. After the last trip, operators lowered the modulator voltage.

Monday May 31
The midnight shift began with Operations monitoring store 3547, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

Operations terminated store 3547 at 1:33 AM after its luminosity dipped below 10E30.

Operations prepared the Collider Detector at Fermilab for access at 1:45 AM. They came out of access at 3:40 AM.

Operations put the TeV into shot setup at 4:25 AM and at 6:31 AM established store 3549 with an initial luminosity of 51.85E30. Pbar resumed stacking.

Plans for Today
The plan for today is to continue with stacking and storing.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 6/1/2004   email Fermilab